# IODeep: an IOD for the introduction of deep learning in the DICOM standard

### 11 Jan 2024 v3

@@ 1 Introduction
In view of the previous considerations, this work provides a complete solution to the problem formulated in Statement 1, and it also represents an enabling technology for answering both Statement 2 and Statement 3. The system proposed in our work leverages the well known DICOM (Digital Imaging and Communications in Medicine)https://www.dicomstandard.org/ standard to achieve the solution, by integrating in the normal DICOM communication between a PACS server and a viewer, the storage and retrieval of a DNN trained purposely on a data set that is related to a particular acquisition modality, anatomical region, and disease. We achieve our stated goal through the creation of a proper Information Object Definition (IOD) that we called IODeep. IODeep contains both the architecture and the weights of a neural network purposely trained on a medical image data set; in our approach, several DNNs are stored in the PACS server, each in a separate IODeep instance, and a novel algorithm is proposed to select the best matching DNN for the images under investigation, based on the information mentioned above that IODeep stores in some proper DICOM tags. Network selection is transparent for the physician who only asks the system to start AI based ROI proposal. The network is used to perform client-side prediction of the Regions of Interest (ROI) on the slice that the physician is currently looking at. On the PACS viewer side, ROIs are visually proposed to the physician who can explicitly validate them as relevant for the diagnosis. The PACS client embeds the set of validated ROIs in an RT Structure Set IOD, which is stored on the server. The entire workflow is conducted in accordance with the DICOM protocol, preserving the security required by the standard, and no modification is required on the PACS server side. The main contributions of the paper are the following:
@@ 2 Theoretical Background
CNN-based networks are often used as a backbone for architectures trained to solve region proposal tasks. In fact, in contrast to semantic segmentation that needs pixel precision, region proposal architectures return predictions about ROIs, i.e. bounding boxes around the truly interesting pattern. Providing loose information about the region to be investigated is the best option for the diagnostic purpose because the physician is not hurried in precise pixel level analysis; rather she or he is simply alerted that some interesting thing may be present in a certain zone of the image. One of the main disadvantages of region proposal architectures is the high computational cost deriving from the increasing number of trainable parameters. Punn and Agarwal [22] in their 2022 survey analysed the state-of-the-art of all U-Net variants used in region proposal tasks; the survey’s emphasis is on the efficiency of the various variants in tricky tasks such as the diagnosis of severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2). Zhang et al. in 2023 [37] propose the MDU-Net, an architecture consisting of a multi-scale dense connections (MDC) encoder and a U-shaped decoder for biomedical image segmentation. The multi-scale dense connections, allowing shorter connections between layers close to the input and output, enable a much deeper U-net that reduces potential overfitting and further improves segmentation performance. Finally, Song et al. [30]. in their work propose the Outlined Attention U-Network, (OAU-Net) based on a bypass branching strategy to solve biomedical image segmentation tasks, capable of detecting surface and deep features. The particularity of this architecture lies in the encoders that are based on residual convolution.
@@ 3 Motivation of the work
In [27] the PyRaDiSe package has been developed, which goes in the direction of a tight integration. PyRaDiSe is an open-source Python package which is independent of DL frameworks, and addresses the issue of artifacts caused by 2D reconstruction as it provides a framework for developing auto-segmentation solutions that can directly operate on DICOM data. Authors claim that PyRaDiSe helps to bridge the gap between data science and clinical radiotherapy by facilitating the implementation of deep learning segmentation models in clinical research practice. Actually, PyRaDiSe has the same research objective as IODeep, and the authors leverage DICOM RT Structured Sets (RTSS) to allow data conversion from DICOM to other image formats thus enabling easy auto-segmentation routines. It is well known in the DICOM related literature [3] that DICOM RTSS are a suitable place to store AI related information like the labels for models training, and also the framework we devised for IODeep makes use of DICOM RTSS to store ROI contours. Differently from PyRaDiSe, in IODeep there is no need to code the segmentation solution. IODeep provides the physician with predictions about relevant ROIs in a completely transparent way.
@@ 4 IODeep architecture
PhotometricInterpretation (0028, 0004) CS RGB or MONOCHROME
@@
Modality (0008, 0060) CS e.g. MR; CT; PT
BodyPartExamined (0018, 0015) CS e.g. BREAST, ABDOMEN, CHEST
@@
Obviously, IODeep is a non-Patient IOD, and it does not contain any patient information. Nevertheless, all the mandatory tags belonging to the “Patient” module (not reported in table 1) along with the other modules composing IODeep, were included in its structure but they were left empty, as stated by the standard. The StudyInstanceUID in the “General Study” module and the SeriesInstanceUID in the “General Series” module are a mandatory tag, but they are not useful in the proposed architecture, and we decided to set them to the same value as the DnnUID tag.
@@
In our implementation the DnnArchitecture tag contains a JSON document whose structure has been developed purposely by some of the authors along with the related parser, which allows to instantiate the same abstract model seamlessly in TensorFlow or PyTorch. Figure 2 shows an example of the JSON describing the architecture showed on the right [18]. As it can be seen, the key input_shape contains the information about the shape of the input tensor, which will be used to check if reshaping is needed for the slice to be processed. The key architecture shows the type and sequential order of the layers that make up the neural network, while the keys convolution, kernel, and so on report the parameters to be set to implement the convolution properly for each layer. In the example a simple sequential model is reported, but our parser deals with all the layer types both in TensorFlow and PyTorch. As an example, our format supports the key skip_connection to build Unet models that are typical architectures for medical imaging.
Figure 2: Representation of a neural network in json format.
In the proposed implementation, we used a PyTorch back-end so the tag DnnWeights contains the URI pointing to a file with .pth extension that is the binary format used by Torch, storing the weights as a serialized Python object.
@@ 4.1 The ROI prediction workflow
In case of successful match, the SOP UID of the IODeep is returned by the algorithm. If no IODeep instance satisfies requirements for network prediction, a warning message will be returned blocking the work pipeline. The returned SOP UID will be used to query the PACS server for obtaining both the network architecture and the URI pointing to the weights file. In turn, the JSON document in the DnnArchitecture tag will be parsed to instantiate an object suitable for building the network instance in the DNN back-end. After the the input tensor shape check, actual prediction will be made. The entire ROI prediction workflow is reported in the Algorithm 2.
@@
Figure 4: The virewer interface for ROI validation. (a) Predicted ROIs are displayed and outlined in red. (b) Validated ROIs are outlined in green.
@@
The RT Structure Set will be uploaded to the PACS server to be reused should the physician wish to retrieve ROI information.
@@ 4.2 Comparison with the DICOM WG-23 proposal
Our proposal moves from a monolithic implementation perspective where information is fully specified according to the DICOM Information Architecture, and only a TCP/IP socket connection is needed to reach the DNN back-end that runs on the same network of the PACS. Apart from the privacy and security concerns already mentioned, we adopted this solution for efficiency reasons. In fact, even if a services ecosystem can be a scalable solution with very low implementation effort on the PACS side, it is tailored on an inference scenario, and it is not well suited to perform training or fine-tuning. Radiology wards are in general focused to particular diseases due to their location in the territory, and the presence of precise medical specializations. As a consequence, fine tuning a trained network by means of the peculiar data generated in a particular PACS infrastructure, appears a reasonable extension of using DNNs in Medical Imaging. Making a data set available at the DNN back-end for training a model is a bandwith consuming task, and a service implementation is not well suited for this. On the other hand, our solution requires only a data transfer on the internal network of the ward, and this is not so resource consuming. IODeep can be used explicitly also in this context: a PACS client devoted to back-end administration could select a data set made by several slice series, and start a network selection procedure, that is the same we described before, followed by the training in the back-end.
@@
Nevertheless, we rearranged our implementation can be easily rearranged according to the indications of the WG-23 work item. The software architecture of this solution is reported in figure 5. In this scenario, Algorithm 1 has been redesigned to provide a service discovery REST API (Application Programming Interface) and it is available at a known endpoint on the service infrastructure. Such an URI is the only thing the PACS platform has to know, and IODeep stores the information used to call the API at the service discovery endpoint. This interaction provides the platform with the endpoint of the selected DNN service. The PACS platform calls the DNN service API at the service endpoint passing the slice, and the information for checking the tensor shape stored in the selected IODeep. Finally, the selected DNN service runs the Algorithm 2. This service architecture does not need to store network information in the IODeep architecture as we can assume that all the networks are already instanced in the DNN back-end. In our implementation the PACS server, the client, and each REST service run in separate Docker containers that communicate throughusing a virtual LAN defined through the YAML configuration file. The actual containers can be deployed on either a single machine or a cluster. We choose explicitly this kind of infrastructure because we strongly believe that effective extension of the service architecture to the training/fine-tuning scenario can be achieved only in a LAN context where the concerns related to the connection bandwidth are negligible.
@@ 5 Conclusion
We strongly believe that the key for using Artificial Intelligence in the everyday diagnostic activity for medical images is related to standardization. Our approach provides a complete answer to the first statement in Section LABEL:sec:intro, and can be reviewed as a technological starting point for answering the other two statements.
@@
Finally, we want to point out that IODeep is also a technological key for explainability of DNN models, and their acceptance in the radiological community. First of all, we claim the need to support the diagnostic process just with a “ROI suggestion” task which is not a precise pixel level prediction that is extremely difficult to be understood by the doctor. Moreover, precise ROI prediction can be achieved only with extensive training on very huge amounts of data. Such a goal can be pursued using Federated Learning (FL) that is a distributed learning paradigm where the model is trained while “moving” it across the data repositories where the data set is split, but data are always kept where they reside without moving them at all. The technology proposed in this paper enables a distributed scenario where IODeeps are shared between the PACS servers of various public and private hospital institutions without violating any rule related to either data security or patients’ privacy, while allowing FL through the data stored within the individual PACS servers.

### 22 Feb 2024 v4

@@ 1 Introduction
In view of the previous considerations, this work provides a complete solution to the problem formulated in Statement 1, and it also represents an enabling technology for answering both Statement 2 and Statement 3. The system proposed in our work leverages the well known DICOM (Digital Imaging and Communications in Medicine)https://www.dicomstandard.org/ standard to achieve the solution, by integrating in the normal DICOM communication between a PACS server and a viewer, the storage and retrieval of a DNN trained purposely on a data set that is related to a particular acquisition modality, anatomical region, and disease. We achieve our stated goal through the creation of a proper Information Object Definition (IOD) that we called IODeep. IODeep contains both the architecture and the weights of a neural network purposely trained on a medical image data set; in our approach, several DNNs are stored in the PACS server, each in a separate IODeep instance, and a novel algorithm is proposed to select the best matching DNN for the images under investigation, based on the information mentioned above that IODeep stores in some proper DICOM tags. Network selection is transparent for the physician who only asks the system to start AI based Regions of Interest (ROI) proposal. The selected network is used to perform client-side prediction of the ROI on the slice that the physician is currently looking at. On the PACS viewer side, ROIs are visually proposed to the physician who can explicitly validate them as relevant for the diagnosis. The PACS client embeds the set of validated ROIs in an RT Structure Set IOD, which is stored on the server. The entire workflow is conducted in accordance with the DICOM protocol, preserving the security required by the standard, and no modification is required on the PACS server side. The main contributions of the paper are the following:
@@ 2 Theoretical Background
CNN-based networks are often used as a backbone for architectures trained to solve region proposal tasks. In fact, in contrast to semantic segmentation that needs pixel precision, region proposal architectures return predictions about ROIs, i.e. bounding boxes around the truly interesting pattern. Punn and Agarwal [22] in their 2022 survey analysed the state-of-the-art of all U-Net variants used in region proposal tasks; the survey’s emphasis is on the efficiency of the various variants in tricky tasks such as the diagnosis of severe acute respiratory syndrome coronavirus 2 (SARS-CoV-2). Zhang et al. in 2023 [39] propose the MDU-Net, an architecture consisting of a multi-scale dense connections (MDC) encoder and a U-shaped decoder for biomedical image segmentation. The multi-scale dense connections, allowing shorter connections between layers close to the input and output, enable a much deeper U-net that reduces potential overfitting and further improves segmentation performance. Finally, Song et al. [31] in their work propose the Outlined Attention U-Network, (OAU-Net) based on a bypass branching strategy to solve biomedical image segmentation tasks, capable of detecting surface and deep features. The particularity of this architecture lies in the encoders that are based on residual convolution.
One of the main disadvantages of region proposal architectures is the high computational cost deriving from the increasing number of trainable parameters. In very recent years, Weakly Supervised Object Location (WSOL) and Weakly Supervised Semantic Segmentation (WSSS) have gained relevance. Weakly supervised techniques do not need explicit explicit pixel-wise labeling to be trained. WSOL [38] is the task of learning how to place a bounding box or a loose ROI around a relevant object in the image, according to the label provided at image level. Many kinds of WSSS [30] are reported in the literature. The most common task in this respect is learning explicit (pixel-wise) segmentation of the object starting from either an image-wise or a box-wise label. Some recent WSSS approaches leverage the Class Activation Maps (CAM) [40]. CAM is a well known XAI approach that generates a map for each class to describe which features were most activated for predicting the class itself. In general, CAM based WSSS use the maps as “seeds” for predicting a first ROI on the object to be segmented. Then the seed ROI is used as a weak mask to train the proper semantic segmentation model [18].
ROI identification through weakly supervised techniques is the best option for the diagnostic purpose because the physician is not hurried in precise pixel level analysis; rather she or he is simply alerted that some interesting thing may be present in a certain zone of the image. Moreover, the model itself does not need to be extremely precise, and this is an advantage in the medical imaging scenario where few training data are at disposal. In view of the previous considerations we decided to design our AI-supported diagnostic scenario as a ROI proposal one. The proposed workflow is agnostic with respect to the use of either bounding boxes or generic ROI contour because, in any case, the vertexes of the ROI can be properly stored in a DICOM RT Structure Set. Moreover, there is no difference in using IODeep to describe a network that learns bounding boxes or a model trained to predict ROIs. Finally, our workflow prescribes that each ROI/bounding box has to be validated manually by the physician; thus we ensure clinical explainability of the whole process.
@@ 3 Motivation of the work
In [Rufenacht__2023] the PyRaDiSe package has been developed, which goes in the direction of a tight integration. PyRaDiSe is an open-source Python package which is independent of DL frameworks, and addresses the issue of artifacts caused by 2D reconstruction as it provides a framework for developing auto-segmentation solutions that can directly operate on DICOM data. Authors claim that PyRaDiSe helps to bridge the gap between data science and clinical radiotherapy by facilitating the implementation of deep learning segmentation models in clinical research practice. Actually, PyRaDiSe has the same research objective as IODeep, and the authors leverage DICOM RT Structure Sets (RTSS) to allow data conversion from DICOM to other image formats thus enabling easy auto-segmentation routines. It is well known in the DICOM related literature [3] that DICOM RTSS are a suitable place to store AI related information like the labels for models training, and also the framework we devised for IODeep makes use of DICOM RTSS to store ROI contours. Differently from PyRaDiSe, in IODeep there is no need to code the segmentation solution. IODeep provides the physician with predictions about relevant ROIs in a completely transparent way.
@@ 4 IODeep architecture
PhotometricInterpretation (0028, 0004) CS
RGB, MONOCHROME1…
(DICOM standard Section C.7.6.3.1.2)
@@
Modality (0008, 0060) CS
CT, MR, PT …
(DICOM standard Section C.7.3.1.1.1)
BodyPartExamined (0018, 0015) CS
BREAST, ABDOMEN, CHEST …
(DICOM Part 16: Content Mapping Resource)
@@
As required by the DICOM WG-23, IODeep is a non-Patient IOD, and it does not contain any patient information. Nevertheless, all the mandatory tags belonging to the “Patient” module (not reported in table 1) along with the other modules composing IODeep, were included in its structure but they were left empty, as stated by the standard. The StudyInstanceUID in the “General Study” module and the SeriesInstanceUID in the “General Series” module are a mandatory tag, but they are not useful in the proposed architecture, and we decided to set them to the same value as the DnnUID tag.
@@
In our implementation the DnnArchitecture tag contains a JSON document whose structure has been developed purposely by some of the authors along with the related parser, which allows to instantiate the same abstract model seamlessly in TensorFlow or PyTorch. The parser executable code is freely available in a suitable repositoryhttps://github.com/CHILab1/DNN-parser linked by the project’s one. The repository contains also the command line specifications to invoke the parser, the complete documentation for writing user-defined JSON descriptions of a network architecture, and the JSON code for building the example U-net [25] used in our implementation. The JSON code is showed also in figure 2 along with the U-net architecture. As it can be seen, the key input_shape contains the information about the shape of the input tensor, which will be used to check if reshaping is needed for the slice to be processed. The key architecture shows the type and sequential order of the layers that make up the neural network, while the keys convolution, kernel, and so on report the parameters to be set to implement the convolution properly for each layer. In the example a simple sequential model is reported, but our parser deals with all the layer types both in TensorFlow and PyTorch. As an example, our format supports the key skip_connection to build U-net models that are typical architectures for medical imaging. A detailed description of the implemented JSON keys, and the executable of the CLI version of the parser are available in the project repository.
Figure 2: Representation of the U-net neural network used in our example, and its JSON description.
In the proposed implementation, we used a PyTorch back-end so the tag DnnWeights contains the URI pointing to a file with .pth extension that is the binary format used by Torch, storing the weights as a serialized Python object. In Tensorfllow we adopted the tf save format to allow the implementation of custom layers. We choose to save separately the weights and the architecture to allow simple creation of the network by the user, who does not need to have strong coding skills. When running the parser, either a torch or a tensorflow environment is set according to the value passed in the --backend command line parameter. The environment is used just to instantiate a specific model that will be saved using the path specified in the --name command line parameter. In this way, a network developer has no need to instantiate a specific model class inherited from either torch.nn.Module or tensorflow.keras.Model. The dcm file containing the IODeep used to instantiate our U-net example network is available in the project repository.
@@ 4.1 The ROI prediction workflow
In case of successful match, the SOP UID of the IODeep is returned by the algorithm. If no IODeep instance satisfies requirements for network prediction, a warning message will be returned blocking the work pipeline. The returned SOP UID will be used to query the PACS server for obtaining both the network architecture and the URI pointing to the weights file. In turn, the JSON document in the DnnArchitecture tag will be parsed to create and save the network instance using the DNN back-end, which is selected according to the value of the backend key. The parser itself is invoked by the PACS client using a system call. After the the input tensor shape check, actual prediction will be made. The entire ROI prediction workflow is reported in the Algorithm 2.
@@
Figure 4: The viewer interface for ROI validation. (a) Predicted ROIs are displayed and outlined in red. (b) Validated ROIs are outlined in green.
@@
The RT Structure Set will be uploaded to the PACS server to be reused should the physician wish to retrieve ROI information. The DICOM file containing the RT Structure Set generated from the ROI displayed in figure 4 is available in the project repository. As already pointed out in section LABEL:sec:back, even if we used an example U-net trained at predicting ROIs, IODeep can be used equally to instantiate both networks that deal with ROIs and models trained at devising bounding boxes. In any case the DICOM RT Structure Set will store the vertex list used to render the poly-line contour in the PACS viewer.
@@ 4.2 Comparison with the DICOM WG-23 proposal
Our proposal moves from a monolithic implementation perspective where information is fully specified according to the DICOM Information Architecture, and only a TCP/IP socket connection is needed to reach the DNN back-end that runs on the same network of the PACS. Apart from the privacy and security concerns already mentioned, we adopted this solution for efficiency reasons. The services ecosystem proposed by the WG-23 is a scalable solution with very low implementation effort on the PACS side, but it is tailored on an inference scenario, and it is not well suited to perform training or fine-tuning. In general, Radiology wards are focused to particular diseases due to their location in the territory, and the presence of precise medical specializations. As a consequence, fine tuning a trained network by means of the peculiar data generated in a particular PACS infrastructure, appears a reasonable extension of using DNNs in Medical Imaging. Making a data set available at the DNN back-end for training a model is a bandwidth consuming task. A service implementation is not well suited for training and/or fine tuning because, in general, a service infrastructure is intended to be hosted in a remote cloud with respect to the applications that require a service. In the model training scenario, this leads to both high latency and privacy issues. On the other hand, our solution requires only a data transfer on the internal network of the ward, and this is not so resource consuming. IODeep can be used explicitly also in this context: a PACS client devoted to back-end administration could select a data set made by several slice series, and start a network selection procedure, that is the same we described before, followed by the training in the back-end.
@@
Nevertheless, we rearranged our implementation according to the indications of the WG-23 work item. The software architecture of this solution is reported in figure 5. In this scenario, Algorithm 1 has been redesigned to provide a service discovery REST API (Application Programming Interface) and it is available at a known endpoint on the service infrastructure. Such an URI is the only thing the PACS platform has to know, and IODeep stores the information used to call the API at the service discovery endpoint. This interaction provides the platform with the endpoint of the selected DNN service. The PACS platform calls the DNN service API at the service endpoint passing the slice, and the information for checking the tensor shape stored in the selected IODeep. Finally, the selected DNN service runs the Algorithm 2. This service architecture does not need to store network information in the IODeep architecture as we can assume that all the networks are already instanced in the DNN back-end. In our implementation the PACS server, the client, and each REST service run in separate Docker containers that communicate throughusing a virtual LAN defined through the YAML configuration file. The actual containers can be deployed on either a single machine or a cluster. We choose explicitly this kind of infrastructure because we strongly believe that effective extension of the service architecture to the training/fine-tuning scenario can be achieved only in a LAN context where the concerns related to the connection bandwidth are negligible.
@@ 5 Conclusion
We strongly believe that the key for using Artificial Intelligence in the everyday diagnostic activity for medical images is related to standardization. Our approach provides a complete answer to the first statement in Section LABEL:sec:intro, and it can be reviewed as a technological starting point for answering the other two statements.
@@
Many DNNs for Medical Imaging work on multi-modal data: this is the case of registered CT-PET volumes, or the very recent generative AIs that integrate both textual and visual information. The current implementation of IODeep does not support multi-modal input data, but this is not a strong limitation for two reasons. First, the general two-stage approach based on both DNN selection and ROI prediction does not depend on the particular input data. Second, the structure of IODeep in Table 1 can be easily extended to utilise multi-modal and/or textual data. In this respect, one can devise a “Study IODeep” that is the proper IOD containing the tags related to both the architecture and the weights of the DNN. Moreover, different “Series IODeep” modules will contain the information related to each kind of input tensor, which can be either a visual or a textual one. Each “Series IODeep” module will reference the DnnUID of the “Study IODeep” to allow proper retrieval of all the information needed to select the multi-modal network, and to check the shape of all the input tensors. We are now working actively on this issue.
Finally, we want to point out that IODeep is also a technological key for explainability of DNN models, and their acceptance in the radiological community. First of all, we claim the need to support the diagnostic process just with a “ROI suggestion” task. A model suggesting a loose ROI does not make a precise pixel level segmentation. In fact, validating an exact segmentation represents a huge cognitive load for the doctor. Moreover, precise ROI prediction can be achieved only with extensive training on very huge amounts of data. Such a goal can be pursued using Federated Learning (FL) that is a distributed learning paradigm where the model is trained while “moving” it across the data repositories where the data set is split, but data are always kept where they reside without moving them at all. The technology proposed in this paper enables a distributed scenario where IODeeps are shared between the PACS servers of various public and private hospital institutions without violating any rule related to either data security or patients’ privacy, while allowing FL through the data stored within the individual PACS servers.